# Designed to Spread: Generative Approaches to Enhance Information Diffusion

### v1

@@ Introduction
Social media has reshaped how people communicate, access information, and form connections. Platforms such as X, Facebook, and TikTok have become powerful channels for news dissemination, public discourse, and cultural influence. Information spreads via diffusion-like processes along user networks, motivating extensive research on diffusion mechanisms (bakshy2012role; guille2013information; weng2013virality).
Extensive research has examined information diffusion in social networks, including modeling diffusion dynamics (granovetter1978threshold; pastor2001epidemic; kempe2003maximizing), identifying influential seed users (chen2009efficient), and limiting spread via edge removal (tong2012gelling). These approaches typically depend on detailed network structure, which is often unavailable or incomplete in real-world settings. Meanwhile, how content is crafted also significantly affects its diffusion, especially for specific audiences. Yet, content-level optimization remains underexplored. This highlights the need for automated methods to generate audience-aware, diffusion-oriented content without relying on explicit network information.
The rapid development of AIGC techniques offers promising potential to fulfill the above needs. While these methods excel at generating high-quality text and image content (chen2023pixartalphafasttrainingdiffusion; podell2023sdxlimprovinglatentdiffusion), optimizing content specifically for maximizing information diffusion remains challenging due to three key issues: (1) diffusion impact is difficult to measure at the content level, especially when network topology is unavailable or incomplete; (2) preserving the original message’s intent while generating user-preferred representations is non-trivial, as misalignment may cause misunderstanding or even unintentional misinformation; and (3) designing a unified framework that integrates diffusion feedback into multimodal content generation remains an open challenge.
@@
Social media analysis explores the complex interplay of factors that shape how information is created, interpreted, and propagated online. Early studies primarily adopted descriptive approaches to reveal observable trends (cheng2014can; de2021no; etta2023characterizing; flamino2023political), drawing insights from three main perspectives.
The first focuses on content virality. For example, Vosoughi et al. (vosoughi2018spread) found that false information spreads more rapidly than truthful content on X, largely due to its novelty and emotional appeal. The second examines user engagement behaviors, including posting frequency (benevenuto2009characterizing; spasojevic2015post), reaction latency (hodas2014simple), and cross-platform activity (xu2014quantifying; iamnitchi2023modeling; alipour2024cross). Especially, Shahbaznezhad et al. (shahbaznezhad2021role) further demonstrated that content format and platform-specific algorithms significantly affect user responsiveness, with emotionally charged posts eliciting faster reactions. The third investigates the structural properties of diffusion cascades. Notarmuzi et al. (notarmuzi2022universality) identified universal propagation patterns in early retweet cascades, highlighting the role of network topology and user influence in determining final reach and virality.
@@
Research on information diffusion models seeks to model the content propagation through social networks. Early approaches were inspired by epidemiological processes, such as the Independent Cascade (IC) and Linear Threshold (LT) models; however, these models assume static networks and fixed influence probabilities among different people (goldenberg2001talk; kempe2003maximizing).
To overcome these simplifications, subsequent work has incorporated temporal dynamics, deep learning, and evolving network structures. For example, NetRate (gomez2012inferring) applies survival analysis to infer time-dependent transmission functions from cascade data. Inf-VAE (sankar2020inf) models latent diffusion representations via variational autoencoders, while DeepDiffuse (islam2018deepdiffuse) leverages deep neural networks to predict both the participants and timing of information cascades. More recently, Meng et al. (meng2025spreading) propose a data-driven model that captures the dynamics of social reinforcement and decay in online diffusion, offering a concise formulation for large-scale spread.
Despite these advancements, existing models still exhibit critical limitations: a strong reliance on explicit network structures, sensitivity to seed selection, and a lack of content awareness in modeling (jiang2023retweet). These challenges underscore the need for a content-based diffusion influence indicator that functions independently of network topology.
@@
Early work on content generation used variational autoencoders (VAEs) (kingma2013auto) and generative adversarial networks (GANs) (goodfellow2020generative) to learn latent representations for attribute-controlled editing (radford2016unsupervisedrepresentationlearningdeep; dumoulin2017adversariallylearnedinference). With large-scale pretrained models, generation has become more fluent and controllable. Large language models (LLMs) such as GPT-3 (brown2020language) and T5 (raffel2020exploring) enable fine-grained text editing via prompts. In vision, diffusion models including Stable Diffusion (rombach2022high), SDXL (podell2023sdxlimprovinglatentdiffusion), and PixArt-$\alpha$ (chen2023pixartalphafasttrainingdiffusion) support high-quality image generation and editing. Multimodal systems such as BLIP-Diffusion (li2023blip) and MIGE (Tian_2025) further unify the editing across modalities.
However, most methods focus on general-purpose generation. Although recent work explores engagement-driven generation for social networks (coppolillo2025engagement), audience-specific preferences remain overlooked. Leveraging the adaptability of LLMs and diffusion models, we explore audience-aware content generation to better support information diffusion.
@@
In this paper, we introduce a new task, Diffusion-Oriented Content Generation (DOCG), which aims to generate a semantically faithful variant (text or image) of a given message $c$, optimized to maximize its diffusion impact among a group of target audiences $U$. We also propose a novel modality-free algorithm framework based on reinforcement learning to accomplish this task. The core idea of our framework is to evaluate the diffusion impact and select strategies to edit the original message iteratively based on pretrained generative models such as GPT-4 (for text) or Pixart-$\alpha$ (chen2023pixartalphafasttrainingdiffusion) (for image). By repeating this evaluation–editing loop, the framework progressively enhances the expected diffusion impact, ultimately producing an optimized variant tailored for the group of target audiences.
@@
Pairwise Diffusion Estimator. The pairwise diffusion estimator is implemented as a neural network that predicts the probability $p_{ij}(c)$. The model operates in three stages: feature extraction, projection, and interaction. First, we extract the features $f_{i}$, $f_{j}$, and $f(c)$ corresponding to users $u_{i}$, $u_{j}$, and content $c$, respectively. User features are computed as the average of features from previously created or shared content, while content features are obtained using a pretrained CLIP encoder (clip), which aligns the text and image features in a shared space. To handle long-text inputs, we adopt Long-CLIP (long-clip), a variant of CLIP designed to overcome token length limitations. Then, all input features are projected into a latent space via a shared Multilayer Perceptron (MLP), denoted as $M_{1}$. The combined representation $h$ is obtained by concatenating the projected vectors:
@@
Action. We define an interpretable action vector $a$, where each dimension $a_{i}\in[-1,1]$ (as detailed in Table 1) corresponds to a specific content feature to be modified. The value of $a_{i}$ indicates the degree and direction of modification: $a_{i}\approx+1$ enhances the $i$-th feature, $a_{i}\approx 0$ leaves it unchanged, and $a_{i}\approx-1$ suppresses it. For text content, we adopt the STEPPS framework (pressgrove2018contagious), which includes six dimensions—Social Currency, Triggers, Emotion, Public, Practical Value, and Stories—each empirically linked to increased virality. For image content, the action space comprises eight perceptual dimensions related to visual engagement, such as Colorfulness, Brightness, and Visual Balance. These dimensions are informed by Li and Xie (li2020picture), who identified image characteristics (e.g., color variation, emotional expression, and photographic quality) as key drivers of user engagement on social media.
@@
where $Q(s,a)$ is the expected reward for applying action $a$ to state $s$, and $\pi_{\theta}$ is the policy parameterized by $\theta$. We optimize the policy network using the policy gradient method (sutton1999policy), which updates parameters to maximize the expected cumulative reward:
@@
Procedure & Metric. We conducted a user study on the Prolific platform (prolific), recruiting native English speakers with informed consent. Participants were screened based on their X activity and recent engagement with the target topic. A total of 100 qualified users were selected (76 male, 24 female; ages: 18–25 (11%), 26–50 (68%), 51–60 (21%)). Each participant evaluated 100 tweet pairs and 100 image pairs, each consisting of an original and a revised version of either text or visual content. For each pair, participants selected the version they were more likely to share (via retweet or reply). To mitigate ordering bias, all pairs were presented in randomized order.

### v2

# Designed to Spread: A Generative Approach to Enhance Information Diffusion

@@ Introduction
Social media has reshaped how people communicate, access information, and form connections. Platforms such as X, Facebook, and TikTok have become powerful channels for news dissemination, public discourse, and cultural influence. Information spreads via diffusion-like processes along user networks, motivating extensive research on diffusion mechanisms (Bakshy et al. 2012; Guille et al. 2013; Weng et al. 2013).
Extensive research has examined information diffusion in social networks, including modeling diffusion dynamics (Granovetter 1978; Pastor-Satorras and Vespignani 2001; Kempe et al. 2003), identifying influential seed users (Chen et al. 2009), and limiting spread via edge removal (Tong et al. 2012). These approaches typically depend on detailed network structure, which is often unavailable or incomplete in real-world settings. Meanwhile, how content is crafted also significantly affects its diffusion, especially for specific audiences. Yet, content-level optimization remains underexplored. This highlights the need for automated methods to generate audience-aware, diffusion-oriented content without relying on explicit network information.
The rapid development of AIGC techniques offers promising potential to fulfill the above needs. While these methods excel at generating high-quality text and image content (Chen et al. 2023; Podell et al. 2023), optimizing content specifically for maximizing information diffusion remains challenging due to three key issues: (1) diffusion impact is difficult to measure at the content level, especially when network topology is unavailable or incomplete; (2) preserving the original message’s intent while generating user-preferred representations is non-trivial, as misalignment may cause misunderstanding or even unintentional misinformation; and (3) designing a unified framework that integrates diffusion feedback into multimodal content generation remains an open challenge.
@@
Social media analysis explores the complex interplay of factors that shape how information is created, interpreted, and propagated online. Early studies primarily adopted descriptive approaches to reveal observable trends (Cheng et al. 2014; De Francisci Morales et al. 2021; Etta et al. 2023; Flamino et al. 2023), drawing insights from three main perspectives.
The first focuses on content virality. For example, Vosoughi et al. (Vosoughi et al. 2018) found that false information spreads more rapidly than truthful content on X, largely due to its novelty and emotional appeal. The second examines user engagement behaviors, including posting frequency (Benevenuto et al. 2009; Spasojevic et al. 2015), reaction latency (Hodas and Lerman 2014), and cross-platform activity (Xu et al. 2014; Iamnitchi et al. 2023; Alipour et al. 2024). Especially, Shahbaznezhad et al. (Shahbaznezhad et al. 2021) further demonstrated that content format and platform-specific algorithms significantly affect user responsiveness, with emotionally charged posts eliciting faster reactions. The third investigates the structural properties of diffusion cascades. Notarmuzi et al. (Notarmuzi et al. 2022) identified universal propagation patterns in early retweet cascades, highlighting the role of network topology and user influence in determining final reach and virality.
@@
Research on information diffusion models seeks to model the content propagation through social networks. Early approaches were inspired by epidemiological processes, such as the Independent Cascade (IC) and Linear Threshold (LT) models; however, these models assume static networks and fixed influence probabilities among different people (Goldenberg et al. 2001; Kempe et al. 2003).
To overcome these simplifications, subsequent work has incorporated temporal dynamics, deep learning, and evolving network structures. For example, NetRate (Gomez-Rodriguez et al. 2012) applies survival analysis to infer time-dependent transmission functions from cascade data. Inf-VAE (Sankar et al. 2020) models latent diffusion representations via variational autoencoders, while DeepDiffuse (Islam et al. 2018) leverages deep neural networks to predict both the participants and timing of information cascades. More recently, Meng et al. (Meng et al. 2025) propose a data-driven model that captures the dynamics of social reinforcement and decay in online diffusion, offering a concise formulation for large-scale spread.
Despite these advancements, existing models still exhibit critical limitations: a strong reliance on explicit network structures, sensitivity to seed selection, and a lack of content awareness in modeling (Jiang et al. 2023). These challenges underscore the need for a content-based diffusion influence indicator that functions independently of network topology.
@@
Early work on content generation used variational autoencoders (VAEs) (Kingma et al. 2013) and generative adversarial networks (GANs) (Goodfellow et al. 2020) to learn latent representations for attribute-controlled editing (Radford et al. 2016; Dumoulin et al. 2017). With large-scale pretrained models, generation has become more fluent and controllable. Large language models (LLMs) such as GPT-3 (Brown et al. 2020) and T5 (Raffel et al. 2020) enable fine-grained text editing via prompts. In vision, diffusion models including Stable Diffusion (Rombach et al. 2022), SDXL (Podell et al. 2023), and PixArt-$\alpha$ (Chen et al. 2023) support high-quality image generation and editing. Multimodal systems such as BLIP-Diffusion (Li et al. 2023) and MIGE (Tian et al. 2025) further unify the editing across modalities.
However, most methods focus on general-purpose generation. Although recent work explores engagement-driven generation for social networks (Coppolillo et al. 2025), audience-specific preferences remain overlooked. Leveraging the adaptability of LLMs and diffusion models, we explore audience-aware content generation to better support information diffusion.
@@
In this paper, we introduce a new task, Diffusion-Oriented Content Generation (DOCG), which aims to generate a semantically faithful variant (text or image) of a given message $c$, optimized to maximize its diffusion impact among a group of target audiences $U$. We also propose a novel modality-free algorithm framework based on reinforcement learning to accomplish this task. The core idea of our framework is to evaluate the diffusion impact and select strategies to edit the original message iteratively based on pretrained generative models such as GPT-4 (for text) or Pixart-$\alpha$ (Chen et al. 2023) (for image). By repeating this evaluation–editing loop, the framework progressively enhances the expected diffusion impact, ultimately producing an optimized variant tailored for the group of target audiences.
@@
Pairwise Diffusion Estimator. The pairwise diffusion estimator is implemented as a neural network that predicts the probability $p_{ij}(c)$. The model operates in three stages: feature extraction, projection, and interaction. First, we extract the features $f_{i}$, $f_{j}$, and $f(c)$ corresponding to users $u_{i}$, $u_{j}$, and content $c$, respectively. User features are computed as the average of features from previously created or shared content, while content features are obtained using a pretrained CLIP encoder (Radford et al. 2021), which aligns the text and image features in a shared space. To handle long-text inputs, we adopt Long-CLIP (Zhang et al. 2024), a variant of CLIP designed to overcome token length limitations. Then, all input features are projected into a latent space via a shared Multilayer Perceptron (MLP), denoted as $M_{1}$. The combined representation $h$ is obtained by concatenating the projected vectors:
@@
Action. We define an interpretable action vector $a$, where each dimension $a_{i}\in[-1,1]$ (as detailed in Table 1) corresponds to a specific content feature to be modified. The value of $a_{i}$ indicates the degree and direction of modification: $a_{i}\approx+1$ enhances the $i$-th feature, $a_{i}\approx 0$ leaves it unchanged, and $a_{i}\approx-1$ suppresses it. For text content, we adopt the STEPPS framework (Pressgrove et al. 2018), which includes six dimensions—Social Currency, Triggers, Emotion, Public, Practical Value, and Stories—each empirically linked to increased virality. For image content, the action space comprises eight perceptual dimensions related to visual engagement, such as Colorfulness, Brightness, and Visual Balance. These dimensions are informed by Li and Xie (Li and Xie 2020), who identified image characteristics (e.g., color variation, emotional expression, and photographic quality) as key drivers of user engagement on social media.
@@
where $Q(s,a)$ is the expected reward for applying action $a$ to state $s$, and $\pi_{\theta}$ is the policy parameterized by $\theta$. We optimize the policy network using the policy gradient method (Sutton et al. 1999), which updates parameters to maximize the expected cumulative reward:
@@
Procedure & Metric. We conducted a user study on the Prolific platform (Prolific 2025), recruiting native English speakers with informed consent. Participants were screened based on their X activity and recent engagement with the target topic. A total of 100 qualified users were selected (76 male, 24 female; ages: 18–25 (11%), 26–50 (68%), 51–60 (21%)). Each participant evaluated 100 tweet pairs and 100 image pairs, each consisting of an original and a revised version of either text or visual content. For each pair, participants selected the version they were more likely to share (via retweet or reply). To mitigate ordering bias, all pairs were presented in randomized order.